# Noise-expansion cascade: an origin of randomness of turbulence

### v1

@@
Abstract Turbulence has remained the greatest challenge in fluid mechanics. Especially, randomness is one of the most important characteristics of turbulence, but its origin is still an open question until now. In this paper, by means of several ingeniously designed “clean” numerical experiments based on Navier-Stokes equations for the two-dimensional turbulent Kolmogorov flow, we reveal such a new phenomenon, namely “noise-expansion cascade”, that all micro-level noises/disturbances at different orders of magnitude in initial condition of Navier-Stokes equations enlarge consistently, say, one by one like an inverse cascade, to macro-level, and besides each of them might greatly change macro-level characteristic of turbulence. The “noise-expansion cascade” closely connects randomness of micro-level noise/disturbance and macro-level disorder of turbulence and thus reveals the origin of randomness of turbulence. In addition, it clearly indicates that unavoidable thermal fluctuation must be considered for turbulence, even if it is several orders of magnitude smaller than other external environmental disturbances.
@@ 1 Motivation
“Turbulence is the last great unsolved problem of classical physics”, as pointed out by the Nobel Prize winner Richard Feynman. Especially, the randomness is one of the most important characteristics of turbulence, but its origin is still an open question until now, to the best of our knowledge.
@@
Today it is widely accepted by scientific community that turbulent flows can be described mathematically by the Navier-Stokes (NS) equations. The NS equations are so important and fundamental that they become the fourth millennium problem [1] of Clay Mathematics Institute of Cambridge, Massachusetts. In 1970 Orszag [2] proposed the so-called “direct numerical simulation” (DNS), which numerically solved the NS equations without any turbulence modes. This is a milestone in fluid mechanics, since it opened the times of numerical experiment. DNS has become a useful tool in fundamental research of turbulence [3, 4, 5, 6, 7], and each result given by DNS is regarded as a “clean” benchmark solution, because it is widely believed that its numerical noise should not be enlarged to a macro-level due to the viscosity of fluid. As pointed out by Coleman and Sandberg [7], DNS has “the ability to perform fundamental studies of clean flows unaffected by numerical, modelling and measurement errors” and “the complete control of the initial and boundary conditions, and each term in the governing equations, also leads to profound advantages over laboratory and field studies.”
@@
Let $f({\bf r})+\Delta_{1}({\bf r})+\Delta_{2}({\bf r})$ denote an initial condition of the NS equations, where ${\bf r}$ is a spatial vector, $f({\bf r})$ is a function at macro-level, $\Delta_{1}({\bf r})$ is a small disturbance at a micro-level, for example at the order $10^{-20}$ of magnitude, and $\Delta_{2}({\bf r})$ is an even smaller disturbance at a micro-level, for example at the order $10^{-40}$ of magnitude, respectively. Traditionally, it is believed that the 2nd disturbance $\Delta_{2}({\bf r})$ could be negligible since it is 20 orders of magnitude smaller than the first disturbance $\Delta_{1}({\bf r})$. Is this traditional viewpoint really correct for turbulent flow governed by NS equations? This is a fundamental problem, which, to the best of our knowledge, is still an open question, too.
@@
To answer the above-mentioned two open questions, it is necessary to develop a new numerical algorithm, whose numerical noise must be much smaller than micro-level physical disturbances and artificial numerical noises of other numerical algorithms throughout a long enough interval of time. In 2009 the so-called “clean numerical simulation” (CNS) was proposed by Liao [8] for chaos and turbulence, and since then its computational efficiency has been increased, step by step, several orders of magnitude [9, 10, 11, 12]. Unlike DNS, CNS algorithm uses multiple-precision [13] with a large enough number of significant digits and thus can decrease both of truncation error and round-off error to any required tiny level. Thus, numerical noise of CNS can be rigorously negligible throughout a time interval that is long enough for calculating statistics [9, 14]. Therefore, CNS result is much more accurate than DNS in a finite but long enough interval of time and thus can be used as “clean” benchmark solution to check the validity of DNS. By comparing the DNS results with the corresponding CNS benchmark solutions in details, it was found [15] that spatio-temporal trajectory of the 2D turbulent Kolmogorov flow given by DNS is badly polluted by artificial numerical noise that quickly increases to the same order of magnitude as the exact solution of NS equations. It illustrated that the 2D turbulent Kolmogorov flow [15] is a chaotic system, say, its spatio-temporal trajectory is rather sensitive to small disturbance caused by artificial numerical noise. More importantly, DNS result sometimes might have great deviations in flow type, spatial symmetry of flow field and even statistics from CNS benchmark solution, as illustrated in [15]. Thus, CNS provides us the ability to do “clean” numerical experiments so that we can accurately investigate the evolution and propagation of micro-level noises/disturbances in initial condition of the NS equations for turbulence. It should be emphasized that this object cannot be realized by DNS whose numerical noise quickly increases to the same order of magnitude as true solution, as illustrated in [15].
@@
Note that CNS has been successfully used to attack some rather difficult problems. For example, the number of periodic orbits of the famous three-body problem, which can be traced back to Newton in 1687, has been increased several orders of magnitude by means of CNS [16, 17, 18], because trajectories of three-body system are essentially chaotic but CNS, unlike other numerical methods, can correctly calculate its chaotic trajectory. Since three-body problem is very famous in history, the discoveries of these new periodic orbits has been reported twice by New Scientist [19, 20]. This is a good example to show the great potential of CNS.
@@
In this paper, several ingeniously designed clean numerical experiments based on CNS are done for a two-dimensional (2D) turbulent Kolmogorov flow governed by NS equations with some specially chosen initial conditions. They provide us a rigorous evidence that all noises/disturbances at different orders of magnitude in initial condition of NS equations could enlarge, one by one like an inverse cascade, to a macro-level, and in addition each of them could lead to great change to turbulence. Based on this phenomenon of our clean numerical experiments, a new concept, namely “noise-expansion cascade”, is proposed. The “noise-expansion cascade” closely connects randomness of micro-level noise/disturbance and macro-level disorder of turbulence, and thus reveals an origin of randomness of turbulence. Besides, according to “noise-expansion cascade”, unavoidable thermal fluctuation must be considered for turbulence, even if it is many orders of magnitudes smaller than other external disturbances.
@@ 2 Set up of clean numerical experiment
3 Results of clean numerical experiment
@@ 3.2 Discovery of “noise-expansion cascade” phenomenon
Figure 1: Comparisons of the vorticity field $\omega(x,y)$ of the 2D turbulent Kolmogorov flow governed by (1) and (2) in the case of $n_{K}=16$ and $Re=2000$ given by CNS, subject to either the initial conditions (3) (left, marked by CNS) or (9) (right, marked by CNS), respectively, at different times, i.e. (a)-(b): $t=30$; (c)-(d): $t=35$; (e)-(f): $t=40$; (g)-(h): $t=300$. For details, please see Movie 1.
@@
Figure 3: Comparisons of the vorticity fields $\omega(x,y)$ of the 2D turbulent Kolmogorov flow governed by (1) and (2) in the case of $n_{K}=16$ and $Re=2000$ given by CNS, subject to either the initial conditions (9) (left, marked by CNS) or (10) (right, marked by CNS′′), respectively, at the different times, i.e. (a)-(b): $t=88$; (c)-(d): $t=93$; (e)-(f): $t=98$; (g)-(h): $t=300$. For details, please see Movie 2.
@@ 4 Conclusion and discussion
It was found [14, 15] that spatiotemporal trajectory given by direct numerical simulation (DNS) is quickly polluted by artificial numerical noise badly so that it is impossible to use DNS to accurately and rigorously simulate the evolution and propagation of small disturbances in turbulent flow. Fortunately, clean numerical simulation (CNS) [9, 10, 11, 12] has much smaller artificial numerical noise than DNS and thus provides us the ability to do “clean” numerical experiments for turbulence. In this paper, by means of several ingeniously designed clean numerical experiments based on CNS for a 2D turbulent Kolmogorov flow, it is discovered, for the first time, that all disturbances at different orders of magnitude in initial condition of NS equations quickly enlarge separately, one by one like an inverse cascade, to macro-level, and besides each of them might greatly change characteristics of the turbulent flow, such as the spatial symmetry of the flow and even the statistics, as illustrated in this paper. Based on this interesting phenomenon of our clean numerical experiments, a new concept, namely “noise-expansion cascade”, is proposed. The so-called “noise-expansion cascade” closely connects randomness of micro-level noise/disturbance and macro-level disorder of turbulence, and thus reveals an origin of randomness of turbulence. This answers our first open question mentioned at the beginning of this paper. It should be emphasized that the “noise-expansion cascade” phenomenon cannot be discovered by numerical experiment based on DNS or in a physical laboratory: this illustrates the potential of CNS.
@@
Note that internal thermal fluctuation and external environmental disturbance are unavoidable in practice. So, according to the “noise-expansion cascade”, thermal fluctuation should be considered for turbulence, even if it is much smaller than external environmental disturbance, just like the Landau-Lifshitz-Navier-Stokes equations that includes the influence of thermal fluctuation. This answers the 2nd open question that we mentioned at the beginning of this paper. Note that the influence of thermal fluctuation on turbulence has been reported currently [25, 26].
@@
Due to the “noise-expansion cascade”, many micro-level random factors might have great influences on macro-level disorder of turbulence: the “noise-expansion cascade” should be a bridge connecting micro-level random fluctuation/disturbance and macro-level disorder of turbulence. So, the new concept “noise-expansion cascade” might provide us a scientific reason why “people cannot step twice into the same river” that was pointed out by philosopher Heraclitus. It highly suggests that turbulence should be a kind of unity of micro-level random noise/disturbance, their evolution and propagation, and macro-level disorder, which are closely connected by the “noise-expansion cascade”: this can also well explain why turbulent flows are so difficult to understand.
@@
Note that, at each time-step, DNS algorithm unavoidably contains artificial numerical noise. Thus, due to the “noise-expansion cascade”, the artificial numerical noise of DNS algorithm at each time-step should increase consistently to macro-level: this is exactly the reason why spatio-temporal trajectory of DNS departs quickly from “true” solution.
@@
There are many interesting problems worthy of further investigation in future. For example, it might be possible that artificial numerical noise could be regarded as a kind of physical external disturbance, as long as we could prove that artificial numerical noise could have same influence as thermal fluctuation and/or environmental disturbance to turbulent flow, which unfortunately is still an open question until now. This kind of investigation can reveal the essence of artificial numerical noises to turbulent flow. Besides, thermal fluctuation is random and discontinuous, but it is still an open question whether or not such random and discontinuous disturbance might become continuous when it increases to macro-level due to the “noise-expansion cascade”.
@@
Finally, we would like to emphasize that DNS is a milestone in fluid mechanics, since it opened a new times of numerical experiment and has greatly promoted the progress of turbulence in theories, physical experiments and applications. In essence, CNS can be regarded as an updated direct numerical simulation with rigorously negligible numerical noises in a finite but long enough interval of time.
@@ Acknowledgements
The calculations were performed on “Tian-He New Generation Supercomputer”, National Supercomputer Center in Tianjing, China. This work is partly supported by National Natural Science Foundation of China (Grant No. 12272230), Postdoctoral Fellowship Program of CPSF (Grant No. GZC20231594) and State Key Laboratory of Ocean Engineering.

### v2

@@
Abstract
Turbulence has remained the greatest challenge in fluid mechanics. Especially, randomness is one of the most important characteristics of turbulence, but its origin is still an open question until now. In this paper, by means of several ingeniously designed “clean” numerical experiments based on Navier-Stokes equations for the two-dimensional turbulent Kolmogorov flow, we reveal such a new phenomenon, namely “noise-expansion cascade”, that all micro-level noises/disturbances at different orders of magnitude in initial condition of Navier-Stokes equations enlarge consistently, say, one by one like an inverse cascade, to macro-level, and besides each of them might greatly change macro-level characteristic of turbulence. The “noise-expansion cascade” closely connects randomness of micro-level noise/disturbance and macro-level disorder of turbulence and thus reveals the origin of randomness of turbulence. In addition, it clearly indicates that unavoidable thermal fluctuation must be considered for turbulence, even if it is several orders of magnitude smaller than other external environmental disturbances.
@@ 1 Motivation
“Turbulence is the last great unsolved problem of classical physics”, as pointed out by the Nobel Prize winner Richard Feynman. Especially, the randomness is one of the most important characteristics of turbulence [1], but its origin is still an open question until now, to the best of our knowledge.
@@
Today it is widely accepted by scientific community that turbulent flows can be described mathematically by the Navier-Stokes (NS) equations. The NS equations are so important and fundamental that they become the fourth millennium problem [2] of Clay Mathematics Institute of Cambridge, Massachusetts. In 1970 Orszag [3] proposed the so-called “direct numerical simulation” (DNS), which numerically solved the NS equations without any turbulence modes. This is a milestone in fluid mechanics, since it opened the times of numerical experiment. DNS has become a useful tool in fundamental research of turbulence [4, 5, 6, 7, 8], and each result given by DNS has been regarded as a “clean” benchmark solution, because it is widely believed that numerical noise of DNS should not be enlarged to a macro-level due to the viscosity of fluid. As pointed out by Coleman and Sandberg [8], DNS has “the ability to perform fundamental studies of clean flows unaffected by numerical, modelling and measurement errors” and “the complete control of the initial and boundary conditions, and each term in the governing equations, also leads to profound advantages over laboratory and field studies.”
@@
Let $f({\bf r})+\Delta_{1}({\bf r})+\Delta_{2}({\bf r})$ denote an initial condition of the NS equations, where ${\bf r}$ is a spatial vector, $f({\bf r})$ is a function at macro-level, $\Delta_{1}({\bf r})$ is a small disturbance at a micro-level, for example at the order $10^{-20}$ of magnitude, and $\Delta_{2}({\bf r})$ is an even smaller disturbance at a micro-level, for example at the order $10^{-40}$ of magnitude, respectively. Traditionally, it is widely believed that the 2nd disturbance $\Delta_{2}({\bf r})$ could be negligible since it is 20 orders of magnitude smaller than the first disturbance $\Delta_{1}({\bf r})$. Is this traditional viewpoint really correct for turbulent flow governed by NS equations? This is a fundamental problem, which, to the best of our knowledge, is still an open question, too.
@@
Obviously, the above-mentioned two open questions have close relationships. In order to answer these open questions, it is necessary to develop a new kind of numerical algorithm, whose numerical noise must be much smaller than micro-level physical disturbances and artificial numerical noises of other traditional numerical algorithms throughout a long enough interval of time. In 2009 the so-called “clean numerical simulation” (CNS) was proposed by Liao [9] for chaos and turbulence, and since then its computational efficiency has been increased, step by step, several orders of magnitude [10, 11, 12, 13]. Unlike DNS, CNS algorithm uses multiple-precision [14] with a large enough number of significant digits and thus can decrease both of truncation error and round-off error to any required tiny level. Thus, numerical noise of CNS can be rigorously negligible throughout a time interval that is long enough for calculating statistics [10]. Therefore, CNS result is much more accurate than DNS in a finite but long enough interval of time and thus can be used as “clean” benchmark solution to check the validity of DNS.
For example, it was found [15] that the DNS result of a two-dimensional turbulent Rayleigh-Bénard (RB) convection, which is excited by a thermal fluctuation as initial condition, quickly departs from the corresponding CNS benchmark solution: it is a kind of non-shearing vortical/roll-like convection at the beginning but then quickly turns into a kind of zonal flow, while the CNS benchmark solution always remains the same non-shearing vortical/roll-like convection in a finite but long enough interval of time $t\in[0,T_{c}]$, where $T_{c}$, called “the critical predictable time”, is equal to 500 (i.e. $T_{c}=500$) for the RB convection under consideration. To further confirm this, a 2D turbulent Kolmogorov flow, which is excited by an initial condition with a spatial symmetry, was solved [16] by DNS and CNS, respectively. It was found [16] that the spatiotemporal trajectory of the CNS benchmark solution remains a kind of spatial symmetry throughout the whole interval of time [0,1000], however the spatiotemporal trajectory of the corresponding DNS result is the same at the beginning as the CNS benchmark result but quickly loses the spatial symmetry, clearly indicating that the spatio-temporal trajectory of the 2D turbulent Kolmogorov flow given by DNS is badly polluted by artificial numerical noise that quickly increases to the same order of magnitude as the exact solution of NS equations. It clearly illustrated that the 2D turbulent Kolmogorov flow [16] is a chaotic system, say, its spatio-temporal trajectory is rather sensitive to small disturbance caused by artificial numerical noise. More importantly, DNS result sometimes might have great deviations from CNS benchmark solution not only in flow type and/or spatial symmetry of flow field but also even in statistics, as illustrated in [15, 16]. These two successful applications of CNS illustrated that CNS can indeed provide us the ability to do such kind of “clean” numerical experiments that we can accurately investigate the evolution and propagation of micro-level noises/disturbances in initial condition of the NS equations for turbulence. It should be emphasized that this object cannot be realized by DNS whose numerical noise quickly increases to the same order of magnitude as true solution, as illustrated in [15, 16].
@@
Note that CNS has been successfully used to attack some rather difficult problems in classical mechanics. For example, the number of periodic orbits of the famous three-body problem, which can be traced back to Newton in 1687, has been increased several orders of magnitude by means of CNS [17, 18, 19], because trajectories of three-body system are essentially chaotic, but CNS, unlike other traditional numerical methods, can correctly calculate its chaotic trajectory. Since three-body problem is very famous in history, the discoveries of these new periodic orbits have been reported twice by New Scientist [20, 21]. This is a good example to show the importance to gain accurate trajectory of some complicated dynamic systems.
@@
In this paper, greatly inspired by spatial symmetry of the CNS benchmark solution of the 2D turbulent Kolmogorov flow [16], we ingeniously design several clean numerical experiments based on CNS for a two-dimensional (2D) turbulent Kolmogorov flow governed by NS equations with some specially chosen initial conditions that contain terms at different orders of magnitudes with different spatial symmetry so as to accurately investigate their propagations, evolutions, and also their macro-scale influences on the turbulence. These clean numerical experiments provide us a rigorous evidence that all noises/disturbances at different orders of magnitude in initial condition of NS equations could enlarge, one by one like an inverse cascade, to a macro-level, and in addition each of them could lead to great change to characteristic of turbulence. Based on this interesting phenomenon, a new concept, namely “noise-expansion cascade”, is proposed. The “noise-expansion cascade” closely connects randomness of micro-level noise/disturbance and macro-level disorder of turbulence, and thus reveals an origin of randomness of turbulence. Besides, according to “noise-expansion cascade”, unavoidable thermal fluctuation must be considered for turbulence, even if it is many orders of magnitudes smaller than other external disturbances.
The set up of the clean numerical experiments is described in § 2. The detailed results of the ingeniously designed clean numerical experiments are reported, and especially the phenomenon, called “noise-expansion cascade”, is revealed in § 3. The conclusions and discussions are given in § 4.
@@ 2 Set up of clean numerical experiment
3 Results of clean numerical experiments
@@ 3.2 Discovery of “noise-expansion cascade” phenomenon
Figure 1: Comparisons of the vorticity field $\omega(x,y)$ of the 2D turbulent Kolmogorov flow governed by (1) and (2) in the case of $n_{K}=16$ and $Re=2000$ given by CNS, subject to either the initial conditions (3) (left, marked by CNS) or (9) (right, marked by CNS), respectively, at different times, i.e. (a)-(b): $t=30$; (c)-(d): $t=35$; (e)-(f): $t=40$; (g)-(h): $t=300$.
@@
Figure 3: Comparisons of the vorticity fields $\omega(x,y)$ of the 2D turbulent Kolmogorov flow governed by (1) and (2) in the case of $n_{K}=16$ and $Re=2000$ given by CNS, subject to either the initial conditions (9) (left, marked by CNS) or (10) (right, marked by CNS′′), respectively, at the different times, i.e. (a)-(b): $t=88$; (c)-(d): $t=93$; (e)-(f): $t=98$; (g)-(h): $t=300$.
@@ 4 Conclusion and discussion
It was found [15, 16] that spatiotemporal trajectory given by “direct numerical simulation” (DNS) is quickly polluted by artificial numerical noise badly so that it is impossible to use DNS to accurately and rigorously simulate the evolution and propagation of small disturbances in turbulent flow. Fortunately, “clean numerical simulation” (CNS) [10, 11, 12, 13] has much smaller artificial numerical noise than DNS and thus provides us the ability to do “clean” numerical experiments for turbulence. In this paper, by means of several ingeniously designed clean numerical experiments based on CNS for a 2D turbulent Kolmogorov flow, it is discovered, for the first time, that all disturbances at different orders of magnitude in initial condition of NS equations quickly enlarge separately, one by one like an inverse cascade, to macro-level, and besides each of them might greatly change characteristics of the turbulent flow not only in the spatial symmetry of the flow but also even in the statistics, as illustrated in this paper. Based on this interesting phenomenon, a new concept, namely “noise-expansion cascade”, is proposed. The so-called “noise-expansion cascade” closely connects randomness of micro-level noise/disturbance and macro-level disorder of turbulence, and thus reveals an origin of randomness of turbulence. This answers our 1st open question mentioned at the beginning of this paper. It should be emphasized that the “noise-expansion cascade” as a new fundamental concept of turbulence cannot be discovered by numerical experiment based on DNS or in a physical laboratory: this illustrates the great potential of CNS.
@@
Note that internal thermal fluctuation and external environmental disturbance are unavoidable in practice. So, according to the “noise-expansion cascade”, thermal fluctuation should be considered for turbulence, even if it is much smaller than external environmental disturbance, just like the Landau-Lifshitz-Navier-Stokes equations that includes the influence of thermal fluctuation[26]. This answers the 2nd open question that we mentioned at the beginning of this paper. Note that the influence of thermal fluctuation on turbulence has been reported currently by few researchers [27, 28].
@@
Due to the “noise-expansion cascade”, many micro-level random factors might have great influences on macro-level disorder of turbulence: the “noise-expansion cascade” should be a bridge connecting micro-level random fluctuation/disturbance and macro-level disorder of turbulence. So, the “noise-expansion cascade” as a new concept might provide us a scientific reason why “people cannot step twice into the same river”, which was pointed out by philosopher Heraclitus. It highly suggests that turbulence should be a kind of unity of micro-level random noise/disturbance, their evolution and propagation, and macro-level disorder, which are closely connected through the “noise-expansion cascade”: this can well explain why turbulence is so difficult to understand.
Note that the “noise-expansion cascade” has obvious difference from “energy cascade”, a fundamental concept in turbulence theory. There are two types of energy cascade, say, direct energy cascade with energy transfer from large scale to small scale, and inverse energy cascade with energy transfer from small scale to large scale. However, unlike energy cascade, noise-expansion cascade has only one direction: from micro-scale to large scale. Besides, energy cascade describes spatial transfer of energy, but noise-expansion cascade describes temporal evolution and propagation of disturbances. This also indicates the novelty of noise-expansion cascade as a new fundamental concept of turbulence.
@@
At each time-step, DNS algorithm unavoidably contains artificial numerical noise. Thus, due to noise-expansion cascade, the artificial numerical noise of DNS algorithm at each time-step should increase consistently to macro-level: this is exactly the reason why spatiotemporal trajectory of DNS departs quickly from true solution. This is the reason why noise-expansion cascade cannot be discovered by DNS.
@@
There are many interesting problems worthy of further investigation in future. For example, it might be possible that artificial numerical noise could be regarded as a kind of physical external disturbance, as long as we could prove that artificial numerical noise could have same influences as thermal fluctuation and/or environmental disturbance to turbulent flow, which unfortunately is still an open question until now. This kind of investigations might reveal the essence of artificial numerical noises to turbulent flow. Besides, thermal fluctuation is random and discontinuous, but it is still an open question whether or not such random and discontinuous disturbance might become continuous when it increases to macro-level due to noise-expansion cascade.
@@
Finally, we would like to emphasize that DNS is a milestone in fluid mechanics, since it opened a new times of numerical experiment and has greatly promoted the progress of turbulence in theories, physical experiments and applications. In essence, CNS can be regarded as a general DNS with rigorously negligible numerical noises in a finite but long enough interval of time, which is powerful and helpful to accurately investigate influence of micro-scale physical/artificial disturbances on turbulent flow.
@@ Acknowledgements
The calculations were performed on “Tian-He New Generation Supercomputer”, National Supercomputer Center in Tianjing, China. This work is partly supported by National Natural Science Foundation of China (Grant No. 12272230 and 91752104), Postdoctoral Fellowship Program of CPSF (Grant No. GZC20231594), and State Key Laboratory of Ocean Engineering.